Evaluate: Ꮾ dx 32 -Ꮖ-2
Answer #1

\\ I = \int_{5}^{\infty} \frac{6}{x^2 - x - 2}dx \\ \\ I = \int_{5}^{\infty} \frac{6}{x^2 - 2*(1/2)*x + (1/4) - 1/4 - 2}dx \\ \\ I = \int_{5}^{\infty} \frac{6}{(x- 1/2)^2 - 9/4}dx \\ \\ Substitute: \ \ x - 1/2 = u \\ \\ dx = du \\ \\ I = \int_{5}^{\infty} \frac{6}{u^2 - 9/4}du \\ \\ I = \int_{5}^{\infty} \frac{24}{4u^2 - 9}du \\ \\ I = 24 \int_{5}^{\infty} \frac{1}{(2u - 3)(2u + 3)}du \\ \\ \frac{1}{(2u - 3)(2u + 3)} = \frac{A}{2u - 3} + \frac{B}{2u + 3} \\ \\ 1 = A(2u + 3) + B(2u - 3) \\ \\ When \ \ u = 3/2, \ \ then \ \ A = \frac{1}{6} \\ \\ When \ \ u = -3/2, \ \ then \ \ B = \frac{-1}{6} \\ \\ So, \\ \\ I = 24 \int_{5}^{\infty} [\frac{1}{6(2u - 3)} - \frac{1}{6(2u + 3)}]du \\ \\ I = 4 \int_{5}^{\infty} [\frac{1}{(2u - 3)} - \frac{1}{(2u + 3)}]du \\

\\ I = 4 \int_{5}^{\infty} [\frac{1}{(2u - 3)} - \frac{1}{(2u + 3)}]du \\ \\ I = 4[\frac{ln (2u - 3)}{2} - \frac{ln (2u + 3)}{2}]_{5}^{\infty} \\ \\ I = 4[\frac{ln (2(x - 1/2) - 3)}{2} - \frac{ln (2*(x - 1/2) + 3)}{2}]_{5}^{\infty} \\ \\ I = 4[\frac{ln (2x - 4)}{2} - \frac{ln (2x + 2)}{2}]_{5}^{\infty} \\ \\ I = 2[ln (\frac{(2x - 4)}{(2x + 2)})]_{5}^{\infty} \\ \\ I = 2[0 - ln (\frac{(2*5 - 4)}{(2*5 + 2)})] \\ \\ \lim_{x \rightarrow \infty} ln (\frac{(2x - 4)}{(2x + 2)}) =ln [ \lim_{x \rightarrow \infty} \frac{(2x - 4)}{(2x + 2)}] = ln (1) = 0 \\ \\ I = 2[0 - ln (\frac{1}{2})] \\ \\ I = {\color{Red} 2*ln 2} \\